# A Survey on LLM-Assisted Clinical Trial Recruitment

### v2

@@ 1 Introduction
Clinical trials evaluate the effects of an intervention on human health. Selecting the precise and required size of patient population is crucial for trial completion. According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time Clinical Trials Arena (2012); Williams et al. (2015); Pharmaceutical Technology (2019). While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research Clinical Trials Arena (2022). As electronic health records (EHRs) of patients become more accessible, clinical researchers adopt machine intelligence and develop explainable systems to correctly interpret model predictions Murdoch and Detsky (2013); Payrovnaziri et al. (2020); von Itzstein et al. (2021).
There are methods leveraging LLMs in clinical trial research for cohort retrieval and modelling Fang et al. (2022); Tian et al. (2023); Park et al. (2024); Liu et al. (2025a); Wang et al. (2025), trial design Reinisch et al. (2024); Curran et al. (2024); Bornet et al. (2025); Neehal et al. (2025), trial search White et al. (2023); Rybinski et al. (2020), trial matching Jin et al. (2024); Nievas et al. (2024); Wornow et al. (2025), predicting trial outcomes and duration Reinisch et al. (2024); Yue et al. (2024a, b); Liu et al. (2025c), assessing risk of bias Lai et al. (2024); Ji et al. (2025), and extracting clinical trial results Lee et al. (2024), even as the community is coming up with recommended practices for responsible use of AI in the entire drug development process Geraci et al. (2025).
@@
Figure 1 shows the components in clinical trial recruitment, namely, data sourcing, information extraction, matching, and evaluation. An expert reviews several hundred patients per trial and can end up spending hours on one patient, hence incurring significant costs Penberthy et al. (2012); Ni et al. (2015). Even simple automation using table queries and lexical searches saves between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation Penberthy et al. (2010). In the past, the patient recruitment process has seen relatively low adoption of the pretrained language models He et al. (2020); Harrer (2023); Lu et al. (2024). Generative LLMs serve as knowledge aggregators, and through their reasoning and instruction-following capabilities, they have revived research in the task of trial and patient matching Jin et al. (2024); Nievas et al. (2024); Rybinski et al. (2024); Wornow et al. (2025).
@@
Despite the rapidly evolving landscape of LLM technology, there is no prior work surveying this area. Gueguen et al. (2025) evaluate public trial matching tools and Layne et al. (2025) compare the efficacy of open and proprietary LLM-assisted trial-patient matching in oncology. Systematic reviews in this topic, bound by strict selection criteria and highly specific research question, do not to capture the broad perspectives Kim and Quintana (2022); Idnay et al. (2022); Kantor and Morzy (2024b); Chen et al. (2025) (Appendix A differentiates between systematic reviews and surveys). Although it provides an overview of LLM approaches used in clinical trial matching, Chen et al. (2025) only briefly discusses associated challenges and future work.
@@ 2 Background
Also known as patient recruitment/enrollment/(pre-)screening, it is the process of matching patients (or a cohort) to a clinical trial via its eligibility criteria. Clinical trials have a dual nature, consisting of both universal and trial-specific requirements, making it challenging to design generalized approaches Idnay et al. (2023). This has traditionally resulted in linear trial-centric matching processes (illustrated in orange in Figure 1) with limited scope for interaction and feedback. Such systems generally have an initial data filter on the structured EHRs followed by keyword matches and concept identification Penberthy et al. (2010); Tun et al. (2023) or cohort-specific classifiers Zhang and Demner-Fushman (2017). Ni et al. (2015) represented both trials and patients as feature vectors, thus supporting retrievals for both trial and patient.
@@
The biomedical NLP landscape is shifting from specialized pretrained language models, such as, BioBERT Lee et al. (2020), BioLM Lewis et al. (2020), PubmedBERT Gu et al. (2021), BioGPT Luo et al. (2022), MedCPT Jin et al. (2023), among many others Wang et al. (2023), towards instruction-following and chat-enabled LLMs (commonly termed as generative AI), used as is Nori et al. (2023); Kung et al. (2023) or fine-tuned for domain alignment, such as, Med-PaLM Singhal et al. (2023), Med-Alpaca Han et al. (2023) and LLaVA-Med Li et al. (2023). We point our reader to Thirunavukarasu et al. (2023); Liu et al. (2025b) for further reading. The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, also observed this shift towards generative AI Lu et al. (2024). Among these, the OpenAI Achiam et al. (2023) family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced models Touvron et al. (2023); Bai et al. (2023); Jiang et al. (2023) models. Further, fine-tuning generative LLMs on the biomedical domain offers limited performance gain Dorfner et al. (2024). In fact, Alber et al. (2025) show that models are more prone to propagate medical misinformation encountered during fine-tuning despite performing well on benchmarks. Finally, Bedi et al. (2025) report that real patient data is used in less than 5% of the 519 biomedical studies using LLMs, and fairness, bias, uncertainty and deployment considerations are rarely assessed.
@@
Kantor and Morzy (2024b)’s study on AI adoption rates in eligibility criteria parsing, report very low adoption rates of generative AI, with BERT family of models being most popular and generative models coming into the picture only since 2024. A systematic review of the role of NLP systems in patient recruitment identified only 11 studies Idnay et al. (2022). Heterogeneous outcomes, diverse results, a dependence on small retrospective data and a lack of common standardized benchmarks drive the gap in NLP research and their adoption in real-world settings Idnay et al. (2024); Kantor and Morzy (2024b). A study by Idnay et al. (2023), investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies. Interestingly, when Corbaux et al. (2024) categorize tools for oncological trial matching, they indicate that the automatic methods still fall in the research and development phase, yet to be commercially available.
@@ 3 Methodology
We queried Google Scholar with the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt” and inspected the first ten result pages, between the years 2019 and 2025 (both inclusive). We also used Jin et al. (2024) and Wornow et al. (2025) as seeds and recursively traced their citations to efficiently captured the evolution of the predictive methods in this field. We organized papers that described LLM-based approaches and corpora for clinical trial recruitment into data sourcing (Section 2), information extraction and parsing (Section 5), trial-patient matching (Section 6) and evaluation (Section 7). Via our discussion on critical limitations in Section 8 and on promising directions towards an interactive patient recruitment in Section 9, we provide further contextualize LLMs in healthcare settings and other stages of the drug development process that directly effect on patient recruitment.
@@
Koopman and Zuccon (2016) #patient summaries (22 words avg) = 60
@@ 11 Query reformulation using LLMs (5),
Best scores are aggregated from the Appendix of the proceedings of TREC 2021. No participants, since this was not a challenge. ∗∗Borrowed from Rybinski et al. (2024) as TREC CT 2023 does not have a published track overview.
@@ 4 Data Sourcing: Public Benchmarks
  • 2018 N2C2 Cohort Selection Stubbs et al. (2019) for criterion-level eligibility prediction. [^1] Currently unavailable as of 2024 Nov 6.
  • Koopman and Zuccon (2016) for ranking trials.
@@ 4.1 Trial Data
The TREC CT benchmarks and Koopman and Zuccon (2016) source the ClinicalTrials.gov registry. Meanwhile, the N2C2 benchmark focuses on a single trial. ClinicalTrials.gov is one of the largest online databases of clinical studies submitted by investigators from over 200 countries. As of June 2025, it lists more than 400,000 clinical trials, thousands of which are active. The euclinicaltrials.eu is another online registry comprising over 50,000 trials from the European Union of which, 7000 are active. Every trial comprises a study description, eligibility criteria and study plan among other details.
@@ 4.2 Patient Data
The N2C2 benchmark contains 288 de-identified longitudinal records of patients. The trial ranking benchmarks use up to 75 synthetic patient profiles, comprising keywords defining cohorts Koopman and Zuccon (2016), short textual patient descriptions (TREC CT 2021, 2022) and questionnaire templates (TREC CT 2023). The MIMIC database Johnson et al. (2016, 2020) is the largest anonymized public database of structured and unstructured data of 299,712 patients. While this dataset is popular for training biomedical LLMs (as mentioned in Section 2), it is not used in any of the five benchmarks, possibly due to significant challenges in obtaining eligibility labels at this scale.
@@ 4.3 Annotated Labels
Two medical experts annotated 3,744 criterion-level labels in the N2C2 benchmark. The TREC CT annotations were created by pooling the top-k results from all participating teams. Medical experts manually annotated this pool of trial-patient matches Koopman and Zuccon (2016); Soboroff (2021); Roberts et al. (2022). Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient Roberts et al. (2022). The relevance judgments were more balanced between the three labels Rybinski et al. (2024) in TREC CT 2023. All benchmarks depend on manual annotation from experts, which is time-consuming and challenging to scale Kim and Quintana (2022).
@@ 5 Information Extraction and Parsing
Extraction of medical entities evolved from a combination of rule-based heuristics and feature-based supervised sequence labelling models Kang et al. (2017); Yuan et al. (2019) via embedding-based neural models Khan et al. (2019); Tseo et al. (2020) to transformer-based pretrained models and LLMs Liu et al. (2021c); Zeng et al. (2020); Tian et al. (2021); Li et al. (2022); Murcia et al. (2024); Kantor and Morzy (2024a). Datta et al. (2024) use disease-specific prompting to extract structured information about entities and its attributes from criteria text. Gao et al. (2020); Zhang et al. (2020); Theodorou et al. (2023) use BERT-based embeddings to encode eligibility criteria and patient data. Patient data converted to search queries, via reformulation and expansion using LLMs, are particularly effective in trial retrieval Peikos et al. (2023); Rybinski et al. (2024); Peikos et al. (2024); Jin et al. (2024). Yuan et al. (2019); Tian et al. (2023); Park et al. (2024); Mugambi et al. (2024); Ziletti and D’Ambrosi (2025) use semantic parsing to translate eligibility criteria into logical forms ready for querying patient databases.
@@ 6.1 Formalization
Despite the task being direction-agnostic, there exist two directional approaches to tackle the matching problem, mainly due to data availability. First is the trial-centric approach, taken by a trial investigator, that matches longitudinal patient records to a specific trial. The 2018 N2C2 cohort selection is a trial-centric benchmark with criterion-level predictions. Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a short description of the patient. Koopman and Zuccon (2016) and the TREC CT 2021, 2022, and 2023 are patient-centric benchmarks for ranking trials.
@@ 6.2 LLM-based Approaches
Unlike classical approaches Penberthy et al. (2010); Zhang and Demner-Fushman (2017); Yuan et al. (2019); Tun et al. (2023), LLM-based approaches do not rely on trial-specific heuristics. Table 2 lists the primary differences between the classical and LLM-based approaches (see Appendix C for a full comparison). We group the LLM-based approaches by the granularity of the matches, starting with criterion-level prediction via trial-level prediction to trial ranking. While criterion- and trial-level predictions are direction-agnostic, trial ranking is a patient-centric task. With enough patients, we could evaluate patient ranking, similar to the bidirectional implementation in Ni et al. (2015)’s work, though this has not yet been addressed.
@@
Here, methods utilize the reasoning capability of LLMs to obtain decision rationale and other context data in addition to the eligibility label. Hamer et al. (2023) use 1-shot prompt, where given the patient profile and the eligibility criteria, the LLM first labels each criterion as being applicable to the patient or not, followed by a list of rationales and finally, the eligibility labels for the applicable criteria. Unlu et al. (2024); Beattie et al. (2024) and Wornow et al. (2025) chunk longitudinal patient data and store them as vector embeddings. They prompt LLMs with criteria and relevant patient data chunks under zero-shot setting, with Beattie et al. (2024) providing expert criterion-level strategy in the prompts. Unlu et al. (2024) generate only the decision label from GPT4 and Beattie et al. (2024) and Wornow et al. (2025) generate criterion-level JSON objects, comprising the criteria label, the eligibility label and a rationale among other context. Both work with proprietary OpenAI models (GPT-3.5 and GPT-4) and Wornow et al. (2025) additionally used open-sourced Llama Jiang et al. (2023) and Mixtral models Jiang et al. (2024). Ferber et al. (2024) prompt GPT-4o to generate criterion-level boolean predictions to be reviewed by experts.
@@
Wong et al. (2023) used GPT3.5 and GPT4 to extract and structure eligibility criteria into logical expression to be matched locally with structured patient information. Yuan et al. (2023) use LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility. The trial-level prediction is then computed by asserting Equation 6.1.
@@
Here, methods formulate effective queries, retrieve trials and re-rank them. Query processing techniques involve generating sentence queries from patient descriptions using a fine-tuned T5 model Pradeep et al. (2022) or zero-shot LLM prompts Saeidi et al. (2023); Kusa et al. (2023b), generating patient descriptions from trial data via 1-shot LLM prompts Zhuang et al. (2023), generating no-SQL queries via LLMs Ferber et al. (2024), using LLMs to reformulate and expand queries Rybinski et al. (2024); Peikos et al. (2024); Datta et al. (2025), and using LLMs to extract keywords Jullien et al. (2024); Jin et al. (2024); Nievas et al. (2024). This is followed by retrieval, either using embeddings similarity only Lahiri et al. (2023); Richmond and Deshpande (2023); Saeidi et al. (2023); Ferber et al. (2024); Saeidi (2025) or a multi-stage retrieval with neural re-rankers Zhuang et al. (2023); Rybinski and Karimi (2023); Rybinski et al. (2024); Jin et al. (2024); Datta et al. (2025). Certain approaches re-rank the top trials using GPT models Zhuang et al. (2023); Rybinski et al. (2024); Datta et al. (2025), some prompt LLMs for relevance labels Pradeep et al. (2022), while others prompt LLMs to generate trial- or criterion-level eligibility labels Rybinski et al. (2024); Peikos (2023); Jin et al. (2024); Nievas et al. (2024); Jullien et al. (2024). Criterion-level labels are aggregated using set-based reasoning mechanisms Jullien et al. (2024), variations of the matching Equation 6.1 Jin et al. (2024); Nievas et al. (2024); Saeidi (2025), and by prompting LLMs Jin et al. (2024).

### v3

@@ 1 Introduction
Clinical trials evaluate the effects of an intervention on human health. Selecting the precise and required size of patient population is crucial for trial completion. According to various estimates, more than 50% of aborted clinical trials fail due to low accrual rates, and 80% of all clinical trials do not manage to recruit the required patient cohorts within the allotted time clinicaltrialarena2012; williams2015terminated; pharmaceutical2019. While the trend has steadily declined over the past decade, with the intensive use of technology-aided solutions, efficient patient recruitment remains the most crucial bottleneck in clinical trial research clinicaltrialarena2022. As electronic health records (EHRs) of patients become more accessible, clinical researchers adopt machine intelligence and develop explainable systems to correctly interpret model predictions murdoch2013inevitable; payrovnaziri2020explainable; vonitzstein2021application.
There are methods leveraging LLMs in clinical trial research for cohort retrieval and modelling fang2022combining; tian2023parsing; park2024criteria2query; liu2025neuralcohort; wang2025trialpanorama, trial design reinisch2024ctp; curran2024examining; bornet2025analysis; neehal2025large, trial search white2023clinidigest; rybinski2020clinical, trial matching jin2024matching; nievas2024distilling; wornow2025zero, predicting trial outcomes and duration reinisch2024ctp; yue2024clinicalagent; yue2024trialdura; liu2025autoct, assessing risk of bias lai2024assessing; ji2025robguard, and extracting clinical trial results lee2024seetrials, even as the community is coming up with recommended practices for responsible use of AI in the entire drug development process geraci2025current.
@@
Figure 1 shows the components in clinical trial recruitment, namely, data sourcing, information extraction, matching, and evaluation. An expert reviews several hundred patients per trial and can end up spending hours on one patient, hence incurring significant costs penberthy2012effort; ni2015increasing. Even simple automation using table queries and lexical searches saves between 165 hours to 1,329 hours of reviewing time when compared to manual evaluation penberthy2010automated. In the past, the patient recruitment process has seen relatively low adoption of the pretrained language models he2020clinical; harrer2023attention; lu2024large. Generative LLMs serve as knowledge aggregators, and through their reasoning and instruction-following capabilities, they have revived research in the task of trial and patient matching jin2024matching; nievas2024distilling; rybinski2024learning; wornow2025zero.
@@
Despite the rapidly evolving landscape of LLM technology, there is no prior work surveying this area. gueguen2025prospective evaluate public trial matching tools and layne2025large compare the efficacy of open and proprietary LLM-assisted trial-patient matching in oncology. Systematic reviews in this topic, bound by strict selection criteria and highly specific research question, do not to capture the broad perspectives kim2022review; idnay2022systematic; kantor2024machine; chen2025enhancing (Appendix A differentiates between systematic reviews and surveys). Although it provides an overview of LLM approaches used in clinical trial matching, chen2025enhancing only briefly discusses associated challenges and future work.
@@ 2 Background
Also known as patient recruitment/enrollment/(pre-)screening, it is the process of matching patients (or a cohort) to a clinical trial via its eligibility criteria. Clinical trials have a dual nature, consisting of both universal and trial-specific requirements, making it challenging to design generalized approaches idnay2023clinical. This has traditionally resulted in linear trial-centric matching processes (illustrated in orange in Figure 1) with limited scope for interaction and feedback. Such systems generally have an initial data filter on the structured EHRs followed by keyword matches and concept identification penberthy2010automated; tun2023automatic or cohort-specific classifiers zhang2017automated. ni2015increasing represented both trials and patients as feature vectors, thus supporting retrievals for both trial and patient.
@@
The biomedical NLP landscape is shifting from specialized pretrained language models, such as, BioBERT lee2020biobert, BioLM lewis2020pretrained, PubmedBERT gu2021domain, BioGPT luo2022biogpt, MedCPT jin2023medcpt, among many others wang2023pre, towards instruction-following and chat-enabled LLMs (commonly termed as generative AI), used as is nori2023capabilities; kung2023performance or fine-tuned for domain alignment, such as, Med-PaLM singhal2023large, Med-Alpaca han2023medalpaca and LLaVA-Med li2023llava. We point our reader to thirunavukarasu2023large; liu2025application for further reading. The Journal of American Medical Informatics Association (JAMIA), which published 41 articles on biomedical health and LLMs in a focus issue, also observed this shift towards generative AI lu2024large. Among these, the OpenAI achiam2023gpt family of proprietary models (GPT3.5 and GPT4 being the most common) is used far more often than open-sourced models touvron2023llama; bai2023qwen; jiang2023mistral models. Further, fine-tuning generative LLMs on the biomedical domain offers limited performance gain dorfner2024biomedical. In fact, alber2025medical show that models are more prone to propagate medical misinformation encountered during fine-tuning despite performing well on benchmarks. Finally, bedi2025testing report that real patient data is used in less than 5% of the 519 biomedical studies using LLMs, and fairness, bias, uncertainty and deployment considerations are rarely assessed.
@@
kantor2024machine’s study on AI adoption rates in eligibility criteria parsing, report very low adoption rates of generative AI, with BERT family of models being most popular and generative models coming into the picture only since 2024. A systematic review of the role of NLP systems in patient recruitment identified only 11 studies idnay2022systematic. Heterogeneous outcomes, diverse results, a dependence on small retrospective data and a lack of common standardized benchmarks drive the gap in NLP research and their adoption in real-world settings idnay2024clinical; kantor2024machine. A study by idnay2023clinical, investigating how clinical researchers screened patients, highlights the challenges of universal and domain-specific nature of the eligibility criteria and makes recommendations to build interactive, flexible and transparent recruitment strategies. Interestingly, when corbaux2024patients categorize tools for oncological trial matching, they indicate that the automatic methods still fall in the research and development phase, yet to be commercially available.
@@ 3 Methodology
We queried Google Scholar with the keywords “clinical trial”, “cohort discovery”, “patient recruitment”, “trial recruitment”, “trial matching” in conjunction with “llm”, “language model”, “gpt” and inspected the first ten result pages, between the years 2019 and 2025 (both inclusive). We also used jin2024matching and wornow2025zero as seeds and recursively traced their citations to efficiently captured the evolution of the predictive methods in this field. We organized papers that described LLM-based approaches and corpora for clinical trial recruitment into data sourcing (Section 2), information extraction and parsing (Section 5), trial-patient matching (Section 6) and evaluation (Section 7). Via our discussion on critical limitations in Section 8 and on promising directions towards an interactive patient recruitment in Section 9, we provide further contextualize LLMs in healthcare settings and other stages of the drug development process that directly effect on patient recruitment.
@@
koopman2016test #patient summaries (22 words avg) = 60
@@ 11 Query reformulation using LLMs (5),
Best scores are aggregated from the Appendix of the proceedings of TREC 2021. No participants, since this was not a challenge. ∗∗Borrowed from rybinski2024learning as TREC CT 2023 does not have a published track overview.
@@ 4 Data Sourcing: Public Benchmarks
  • 2018 N2C2 Cohort Selection stubbs2019cohort for criterion-level eligibility prediction. [^1] Currently unavailable as of 2024 Nov 6.
  • koopman2016test for ranking trials.
@@ 4.1 Trial Data
The TREC CT benchmarks and koopman2016test source the ClinicalTrials.gov registry. Meanwhile, the N2C2 benchmark focuses on a single trial. ClinicalTrials.gov is one of the largest online databases of clinical studies submitted by investigators from over 200 countries. As of June 2025, it lists more than 400,000 clinical trials, thousands of which are active. The euclinicaltrials.eu is another online registry comprising over 50,000 trials from the European Union of which, 7000 are active. Every trial comprises a study description, eligibility criteria and study plan among other details.
@@ 4.2 Patient Data
The N2C2 benchmark contains 288 de-identified longitudinal records of patients. The trial ranking benchmarks use up to 75 synthetic patient profiles, comprising keywords defining cohorts koopman2016test, short textual patient descriptions (TREC CT 2021, 2022) and questionnaire templates (TREC CT 2023). The MIMIC database johnson2016mimic; johnson2020mimic is the largest anonymized public database of structured and unstructured data of 299,712 patients. While this dataset is popular for training biomedical LLMs (as mentioned in Section 2), it is not used in any of the five benchmarks, possibly due to significant challenges in obtaining eligibility labels at this scale.
@@ 4.3 Annotated Labels
Two medical experts annotated 3,744 criterion-level labels in the N2C2 benchmark. The TREC CT annotations were created by pooling the top-k results from all participating teams. Medical experts manually annotated this pool of trial-patient matches koopman2016test; soboroff2021overview; roberts2022overview. Out of a total of 35,394 pooled trial-patient matches in the 2022 edition, 11% were judged as eligible, 9% as ineligible and 80% as not relevant with an average of 700 trials judged per patient roberts2022overview. The relevance judgments were more balanced between the three labels rybinski2024learning in TREC CT 2023. All benchmarks depend on manual annotation from experts, which is time-consuming and challenging to scale kim2022review.
@@ 5 Information Extraction and Parsing
Extraction of medical entities evolved from a combination of rule-based heuristics and feature-based supervised sequence labelling models kang2017eliie; yuan2019criteria2query via embedding-based neural models khan2019improving; tseo2020information to transformer-based pretrained models and LLMs liu2021clinical; zeng2020ensemble; tian2021transformer; li2022comparative; murcia2024automating; kantor2024finetuned. datta2024autocriteria use disease-specific prompting to extract structured information about entities and its attributes from criteria text. gao2020compose; zhang2020deepenroll; theodorou2023treement use BERT-based embeddings to encode eligibility criteria and patient data. Patient data converted to search queries, via reformulation and expansion using LLMs, are particularly effective in trial retrieval peikos2023investigating; rybinski2024learning; peikos2024leveraging; jin2024matching. yuan2019criteria2query; tian2023parsing; park2024criteria2query; mugambi2024leveraging; ziletti2025generating use semantic parsing to translate eligibility criteria into logical forms ready for querying patient databases.
@@ 6.1 Formalization
Despite the task being direction-agnostic, there exist two directional approaches to tackle the matching problem, mainly due to data availability. First is the trial-centric approach, taken by a trial investigator, that matches longitudinal patient records to a specific trial. The 2018 N2C2 cohort selection is a trial-centric benchmark with criterion-level predictions. Alternatively, a patient-centric approach, taken by a patient or their healthcare provider, matches relevant trials from a trial registry to a short description of the patient. koopman2016test and the TREC CT 2021, 2022, and 2023 are patient-centric benchmarks for ranking trials.
@@ 6.2 LLM-based Approaches
Unlike classical approaches penberthy2010automated; zhang2017automated; yuan2019criteria2query; tun2023automatic, LLM-based approaches do not rely on trial-specific heuristics. Table 2 lists the primary differences between the classical and LLM-based approaches (see Appendix C for a full comparison). We group the LLM-based approaches by the granularity of the matches, starting with criterion-level prediction via trial-level prediction to trial ranking. While criterion- and trial-level predictions are direction-agnostic, trial ranking is a patient-centric task. With enough patients, we could evaluate patient ranking, similar to the bidirectional implementation in ni2015increasing’s work, though this has not yet been addressed.
@@
Here, methods utilize the reasoning capability of LLMs to obtain decision rationale and other context data in addition to the eligibility label. hamer2023improving use 1-shot prompt, where given the patient profile and the eligibility criteria, the LLM first labels each criterion as being applicable to the patient or not, followed by a list of rationales and finally, the eligibility labels for the applicable criteria. unlu2024retrieval; beattie2024utilizing and wornow2025zero chunk longitudinal patient data and store them as vector embeddings. They prompt LLMs with criteria and relevant patient data chunks under zero-shot setting, with beattie2024utilizing providing expert criterion-level strategy in the prompts. unlu2024retrieval generate only the decision label from GPT4 and beattie2024utilizing and wornow2025zero generate criterion-level JSON objects, comprising the criteria label, the eligibility label and a rationale among other context. Both work with proprietary OpenAI models (GPT-3.5 and GPT-4) and wornow2025zero additionally used open-sourced Llama jiang2023mistral and Mixtral models jiang2024mixtral. ferber2024end prompt GPT-4o to generate criterion-level boolean predictions to be reviewed by experts.
@@
wong2023scaling used GPT3.5 and GPT4 to extract and structure eligibility criteria into logical expression to be matched locally with structured patient information. yuan2023large use LLMs to augment eligibility criteria and pass the BERT-based embeddings of these criteria and patient data through a fully connected classification layer to predict patient-criterion eligibility. The trial-level prediction is then computed by asserting Equation 6.1.
@@
Here, methods formulate effective queries, retrieve trials and re-rank them. Query processing techniques involve generating sentence queries from patient descriptions using a fine-tuned T5 model pradeep2022neural or zero-shot LLM prompts saeidi2023malnis; kusa2023dossier, generating patient descriptions from trial data via 1-shot LLM prompts zhuang2023team, generating no-SQL queries via LLMs ferber2024end, using LLMs to reformulate and expand queries rybinski2024learning; peikos2024leveraging; datta2025patient2trial, and using LLMs to extract keywords jullien2024controlled; jin2024matching; nievas2024distilling. This is followed by retrieval, either using embeddings similarity only lahiri2023tmu; richmond2023leveraging; saeidi2023malnis; ferber2024end; saeidi2025streamlining or a multi-stage retrieval with neural re-rankers zhuang2023team; rybinski2023matching; rybinski2024learning; jin2024matching; datta2025patient2trial. Certain approaches re-rank the top trials using GPT models zhuang2023team; rybinski2024learning; datta2025patient2trial, some prompt LLMs for relevance labels pradeep2022neural, while others prompt LLMs to generate trial- or criterion-level eligibility labels rybinski2024learning; piekos2023unimib; jin2024matching; nievas2024distilling; jullien2024controlled. Criterion-level labels are aggregated using set-based reasoning mechanisms jullien2024controlled, variations of the matching Equation 6.1 jin2024matching; nievas2024distilling; saeidi2025streamlining, and by prompting LLMs jin2024matching.